# Persistent and anti-Persistent Motion in Bounded and Unbounded Space: Resolution of the First-Passage Problem

### v1

@@ V Absorbing Traps and First-Passage Processes
We are interested in studying first-passage processes both when a single target consists of a subset of the $i=1,...,M$ internal degrees of freedom or all of them (the entire site), in the cases where the initial probability is shared between any or all of the internal states in $\bm{n}_{0}$. Since the set of trajectories that originate in state $\bm{n}_{0},m_{i}$ are independent of those starting in $\bm{n}_{0},m_{j}$ and those terminating in $\bm{s},m_{\bm{s}_{i}}$ are independent of those terminating in $\bm{s},m_{\bm{s}_{j}}$ $(i\neq j)$, it suffices to consider them separately and sum the result. Furthermore, when there are multiple targets, the summation is carried out over both the internal states and the site location.
In the presence of an unordered set of absorbing traps $\mathdutchcal{S}=\left\{(\bm{s},m_{\bm{s}})_{i},...,(\bm{s},m_{\bm{s}})_{S}\right\}$ ($S$ is the cardinality of the set), where each element represents a trap in the internal state at $(\bm{s}_{i},m_{\bm{s}_{i}})$, we obtain (see C for the details) the following propagator generating function
$\widetilde{P}^{(a)}_{\bm{n}_{0}}(\bm{n},z)=\widetilde{P}_{\bm{n}_{0}}(\bm{n},z)-\sum_{m=1}^{M}\sum_{j=1}^{M}\sum_{i=1}^{S}\alpha_{m_{\bm{s}_{i}}}\widetilde{P}_{\bm{s},m_{\bm{s}_{i}}}(\bm{n},m,z)\frac{\det[\mathbb{H}^{(i)}(\bm{n}_{0},m_{0_{j}},z)]}{\det[\mathbb{H}(z)]},$ (21)
@@
where $\mathbb{H}(z)_{l,k}=\alpha_{m_{\bm{s}_{k}}}\widetilde{P}_{\bm{s}_{k},m_{\bm{s}_{k}}}(\bm{s}_{l},m_{\bm{s}_{l}},z)$, $\mathbb{H}(z)_{k,k}=\alpha_{m_{\bm{s}_{k}}}\widetilde{P}_{\bm{s}_{k},m_{\bm{s}_{k}}}(\bm{s}_{k},m_{\bm{s}_{k}},z)$ and $\mathbb{H}^{(i)}(\bm{n}_{0},m_{0},z)$ is the same, but with the $i^{\text{th}}$ column replaced with $\alpha_{m_{0_{j}}}\left[\widetilde{P}_{\bm{n}_{0},m_{0_{j}}}(\bm{s}_{1},m_{\bm{s}_{1}},z),...,\widetilde{P}_{\bm{n}_{0},m_{0_{j}}}(\bm{s}_{S},m_{\bm{s}_{S}},z)\right]^{\intercal}$. The notation $\alpha_{m_{0_{j}}}P_{\bm{n}_{0},m_{0_{j}}}(\bm{n},m,t)=\alpha_{m_{0_{j}}}\bm{e}_{m}^{\intercal}\cdot\bm{P}_{\bm{n}_{0}}(\bm{n},t)\cdot\bm{e}_{m_{0_{j}}}$, is the occupation probability that the walker occupies site and state $(\bm{n},m)$ given its initial position was $(\bm{n}_{0},m_{0_{j}})$, which is weighted by $\alpha_{m_{0_{j}}}$.
@@
Depending on the problem Eq. (22) can be simplified to more compact expressions. The simplest case is the one target state for the one-dimensional unbounded CRW. In such a scenario, if the target state is located at site $s>n_{0}$, one may put the target at $(s,1)$ since the walker may only arrive at $s$ from the left, as one can evince by looking at Fig. 1(b). With a single target state, Eq. (22) can be written as
$\widetilde{F}_{\bm{n}_{0},m_{0}}(\bm{s},m_{\bm{s}},z)=\frac{\sum_{j=1}^{M}\alpha_{m_{0_{j}}}\widetilde{P}_{n_{0},m_{0_{j}}}(\bm{s},m_{\bm{s}},z)}{\widetilde{P}_{\bm{s},m_{\bm{s}}}(\bm{s},m_{\bm{s}},z)},$ (24)
@@
$\widetilde{F}_{n_{0}}(s,z)$ $=\frac{\alpha_{1}\!\left(\!\widetilde{P}_{n_{0},1}(s,1,z)\left[\widetilde{P}_{s,2}(s,2,z)\!-\!\widetilde{P}_{s,1}(s,2,z)\right]\!+\!\widetilde{P}_{n_{0},1}(s,2,z)\!\left[\widetilde{P}_{s,1}(s,1,z)\!-\!\widetilde{P}_{s,2}(s,1,z)\right]\!\right)}{\widetilde{P}_{s,1}(s,1,z)\widetilde{P}_{s,2}(s,2,z)-\widetilde{P}_{s,1}(s,2,z)\widetilde{P}_{s,2}(s,1,z)\!}$
@@
where we have lightened the notation with $\widetilde{F}_{n_{0}}(s,z)=\widetilde{F}_{n_{0}}(\{(s,1),(s,2)\},z)$, $\alpha_{m_{0_{1}}}=\alpha_{1}$ and similarly for the second state weighting. The alternative derivation via a generalised renewal equation of Eqs. (24) and (25) can be found in Sec. VII of ref. [41].
@@
The sustained oscillations, which have only been seen in a small number of systems, including quantum walks [71], biased random walks in hexagonal domains [69] and other non-Markov systems [72] are the natural, higher-dimensional analogue of the bimodality seen in the one-dimensional systems, where again, the peaks correspond to the timescales for which the walker moves away from the target as well as those trajectories that miss the target and must travel round the entire domain. Naturally, as the indirect trajectories take over and more and more trajectories reach the targets, the oscillations dampen and decay to zero [69].
@@ VI Summary Statistics and the Mean First-Passage Time
The matrices in Eq. (26) are made up of the MFPT between localised sites $\mathdutchcal{F}_{\bm{n}_{0},m_{0}\to\bm{s}_{i},m_{\bm{s}_{i}}}$. More precisely, $\mathbb{T}_{i,i}=0$, $\mathbb{T}_{i,k}=\mathdutchcal{F}_{\bm{s}_{k},m_{\bm{s}_{k}}\to\bm{s}_{i},m_{\bm{s}_{i}}}$ ($i\neq k$), $\mathbb{T}_{{0}_{i,k}}=\mathbb{T}_{i,k}-(2d)^{-1}\sum_{j=1}^{2d}\mathdutchcal{F}_{\bm{n}_{0},m_{0_{j}}\to\bm{s}_{i},m_{\bm{s}_{i}}}$ and $\mathbb{T}_{1_{i,k}}=\mathbb{T}_{i,k}-1$. To find these quantities one evaluates expressions of the form $\mathdutchcal{F}_{\bm{n}_{0},m_{0}\to\bm{s}_{i},m_{\bm{s}_{i}}}=\frac{\partial}{\partial z}\left.\frac{\widetilde{P}_{\bm{n}_{0},m_{0}}(\bm{s},m_{\bm{s}},z)}{\widetilde{P}_{\bm{s},m_{\bm{s}}}(\bm{s},m_{\bm{s}},z)}\right|_{z=1},$ and we give details and explicit quantities in Sec. VIII of ref. [41]. Performing a summation of Eq. (26) over $\bm{n}_{0}$ yields the global mean first-passage time (GMFPT) [35], which we denote as $\mathdutchcal{G}_{\{\mathdutchcal{S}\}}$. In such a case, one replaces $\mathbb{T}_{0}$ with $\mathbb{G}$ where $\mathbb{G}_{i,j}=\mathbb{T}_{i,j}-(2dN)^{-1}\sum_{\bm{n}_{0}}\sum_{j=1}^{2d}\mathdutchcal{F}_{\bm{n}_{0},m_{0_{j}}\to\bm{s}_{i},m_{\bm{s}_{i}}}$, while $\mathbb{T}$ and $\mathbb{T}_{1}$ are unaffected by the summation. We discuss this quantity further in ref. [41].

### v2

@@ V Absorbing Traps and First-Passage Processes
We are interested in studying first-passage processes both when a single target consists of a subset of the $i=1,...,M$ internal degrees of freedom or all of them (the entire site), in the cases where the initial probability is shared between any or all of the internal states in $\bm{n}_{0}$. Since the set of trajectories that originate in state $\bm{n}_{0},m_{i}$ are independent of those starting in $\bm{n}_{0},m_{j}$ and those terminating in $\bm{s},m_{\bm{s}_{i}}$ are independent of those terminating in $\bm{s},m_{\bm{s}_{j}}$ $(i\neq j)$, it suffices to consider them separately and sum the result. Furthermore, when there are multiple targets, the summation is carried out over both the internal states and the physical site locations.
To illustrate, consider the (unordered) set $\mathdutchcal{S}=\left\{(\bm{s}_{i},m_{\bm{s}_{i}}),...,(\bm{s}_{S},m_{\bm{s}_{S}})\right\}$ comprising of $S$ localised targets. To begin, we write, in arbitrary dimensions, the Master equation for the state-level dynamics
$P(\bm{n},m,t+1)=\sum_{\bm{n}^{\prime}}\sum_{m^{\prime}=1}^{M}$ $\bigg{[}A(\bm{n},m,\bm{n}^{\prime},m^{\prime})P(\bm{n}^{\prime},m^{\prime},t)$
$+\sum_{i=1}^{S}\delta_{\bm{n},\bm{s}_{i}}\delta_{m,m_{\bm{s}_{i}}}(1-\rho_{\bm{s}_{i},m_{\bm{s}_{i}}})A(\bm{s}_{i},m_{\bm{s}_{i}},\bm{n}^{\prime},m^{\prime})P(\bm{n}^{\prime},m^{\prime},t)\bigg{]},$ (21)
where $A(\bm{n},m,\bm{n^{\prime}},m^{\prime})$ is the transition probability from state $\bm{n}^{\prime},m^{\prime}$ to state $\bm{n},m$, and $\rho_{\bm{s}_{i},m_{\bm{s}_{i}}}$, $(0<\rho_{\bm{s}_{i},m_{\bm{s}_{i}}}\leq 1)$ governs the probability of getting absorbed at defect $\bm{s}_{i},m_{\bm{s}_{i}}$ where $\rho_{\bm{s}_{i},m_{\bm{s}_{i}}}=1$ represents perfect trapping efficiency at that site. To proceed, we first consider the $\rho_{\bm{s}_{i},m_{\bm{s}_{i}}}\neq 1$ case.
$\widetilde{P}^{(a)}(\bm{n},m,z)$ $=\!\sum_{\bm{n}_{0}^{\prime}}\sum_{m_{0}^{\prime}=1}^{M}\!\bigg{[}\widetilde{P}_{\bm{n}_{0}^{\prime},m_{0}^{\prime}}(\bm{n},m,z)P(\bm{n},m,0)$
$+\sum_{i=1}^{S}\frac{\rho_{\bm{s}_{i},m_{\bm{s}_{i}}}}{\rho_{\bm{s}_{i},m_{\bm{s}_{i}}}\!-\!1}\widetilde{P}_{\bm{s}_{i},m_{\bm{s}_{i}}}(\bm{n},m,\!z)\!\left[\widetilde{P}^{(a)}(\bm{s}_{i},m_{\bm{s}_{i}},\!z)\!-\!P^{(a)}(\bm{s}_{i},m_{\bm{s}_{i}},0)\right]\!\bigg{]}\!,$ (23)
where we have used $\widetilde{P}_{\bm{n}_{0},m_{0}}(\bm{n},m,z)$ to denote any valid defect free propagator.
$\widetilde{P}^{(a)}_{\bm{n}_{0},m_{0}}(\bm{n},m,z)$ $=\alpha_{m_{0}}\widetilde{P}_{\bm{n}_{0},m_{0}}(\bm{n},m,z)$
$+\sum_{i=1}^{S}\frac{\rho_{\bm{s}_{i},m_{\bm{s}_{i}}}}{\rho_{\bm{s}_{i},m_{\bm{s}_{i}}}-1}\alpha_{m_{\bm{s}_{i}}}\widetilde{P}_{\bm{s}_{i},m_{\bm{s}_{i}}}(\bm{n},m,z)\widetilde{P}^{(a)}_{\bm{n}_{0},m_{0}}(\bm{s}_{i},m_{\bm{s}_{i}},z),$ (24)
@@
where the notation $\alpha_{m_{0_{j}}}P_{\bm{n}_{0},m_{0_{j}}}(\bm{n},m,t)=\alpha_{m_{0_{j}}}\bm{e}_{m}^{\intercal}\cdot\bm{\mathcal{P}}_{\bm{n}_{0}}(\bm{n},t)\cdot\bm{e}_{m_{0_{j}}}$, where $\bm{P}_{\bm{n}_{0}}(\bm{n},t)=\bm{\mathcal{P}}_{\bm{n}_{0}}(\bm{n},t)\cdot\bm{U}_{m_{0}}$, is the probability that the walker occupies site and state $(\bm{n},m)$ given its initial position was $(\bm{n}_{0},m_{0_{j}})$, which is weighted by $\alpha_{m_{0_{j}}}$.
Following the standard defect technique procedure [39] i.e., solving via Cramer’s rule and taking $\rho_{\bm{s}_{i},m_{\bm{s}_{i}}}\to 1$ for all $(\bm{s}_{i},m_{\bm{s}_{i}})$, we find the state-level propagator with absorption as
$\widetilde{P}^{(a)}_{\bm{n}_{0},m_{0}}(\bm{n},m,z)=\alpha_{m_{0}}\widetilde{P}_{\bm{n}_{0},m_{0}}(\bm{n},m,z)-\sum_{i=1}^{S}\alpha_{m_{\bm{s}_{i}}}\widetilde{P}_{\bm{s}_{i},m_{\bm{s}_{i}}}(\bm{n},m,z)\frac{\det[\mathbb{H}^{(i)}(\bm{n}_{0},m_{0},z)]}{\det[\mathbb{H}(z)]},$ (25)
where $\mathbb{H}(z)_{l,k}=\alpha_{m_{\bm{s}_{k}}}\widetilde{P}_{\bm{s}_{k},m_{\bm{s}_{k}}}(\bm{s}_{l},m_{\bm{s}_{l}},z)$, $\mathbb{H}(z)_{k,k}=\alpha_{m_{\bm{s}_{k}}}\widetilde{P}_{\bm{s}_{k},m_{\bm{s}_{k}}}(\bm{s}_{k},m_{\bm{s}_{k}},z)$ and $\mathbb{H}^{(j)}(\bm{n}_{0},m_{0},z)$ is the same but with the $j^{\text{th}}$ column replaced with $\alpha_{m_{0}}\left[\widetilde{P}_{\bm{n}_{0},m_{0}}(\bm{s}_{1},m_{\bm{s}_{1}},z),...,\widetilde{P}_{\bm{n}_{0},m_{0}}(\bm{s}_{S},m_{\bm{s}_{S}},z)\right]^{\intercal}$, i.e, the matricies are comprised of known state-level propagators.
As noted above, the occupation probability over the entire site may be found by performing a summation over all internal states and using the fact that $\widetilde{P}^{(a)}_{\bm{n}_{0}}(\bm{n},z)=\sum_{m=1}^{M}\sum_{j=1}^{M}\widetilde{P}_{\bm{n}_{0},m_{0_{j}}}^{(a)}(\bm{n},m,z)$, and $\widetilde{P}^{(\gamma)}_{\bm{n}_{0}}(\bm{n},z)=\sum_{m=1}^{M}\sum_{j=1}^{M}\alpha_{m_{0_{j}}}\widetilde{P}_{\bm{n}_{0},m_{0_{j}}}^{(\gamma)}(\bm{n},m,z)$, where there is no $\alpha_{m_{0}}$ multiplier on the absorbing case as it is already implemented via the initial condition in Eq. (24), one may perform the double summation over both sides of Eq. (25) to arrive at
$\widetilde{P}^{(a)}_{\bm{n}_{0}}(\bm{n},z)=\widetilde{P}_{\bm{n}_{0}}(\bm{n},z)-\sum_{m=1}^{M}\sum_{j=1}^{M}\sum_{i=1}^{S}\alpha_{m_{\bm{s}_{i}}}\widetilde{P}_{\bm{s},m_{\bm{s}_{i}}}(\bm{n},m,z)\frac{\det[\mathbb{H}^{(i)}(\bm{n}_{0},m_{0_{j}},z)]}{\det[\mathbb{H}(z)]}.$ (26)
@@
Depending on the problem Eq. (27) can be simplified to more compact expressions. The simplest case is that of one target state, which is pertinent to the one-dimensional unbounded CRW. In such a scenario, if the target state is located at site $s>n_{0}$, one may put the target at $(s,1)$ since the walker may only arrive at $s$ from the left, as one can evince by looking at Fig. 2(b). With a single target state, Eq. (27) can be written as
$\widetilde{F}_{\bm{n}_{0}}(\bm{s},m_{\bm{s}},z)=\frac{\sum_{j=1}^{M}\alpha_{m_{0_{j}}}\widetilde{P}_{n_{0},m_{0_{j}}}(\bm{s},m_{\bm{s}},z)}{\widetilde{P}_{\bm{s},m_{\bm{s}}}(\bm{s},m_{\bm{s}},z)},$ (29)
@@
$\widetilde{F}_{n_{0}}(s,z)=$
$\frac{\alpha_{1}\!\left(\!\widetilde{P}_{n_{0},1}(s,1,z)\left[\widetilde{P}_{s,2}(s,2,z)\!-\!\widetilde{P}_{s,1}(s,2,z)\right]\!+\!\widetilde{P}_{n_{0},1}(s,2,z)\!\left[\widetilde{P}_{s,1}(s,1,z)\!-\!\widetilde{P}_{s,2}(s,1,z)\right]\!\right)}{\widetilde{P}_{s,1}(s,1,z)\widetilde{P}_{s,2}(s,2,z)-\widetilde{P}_{s,1}(s,2,z)\widetilde{P}_{s,2}(s,1,z)\!}$
@@
where we have lightened the notation with $\widetilde{F}_{n_{0}}(s,z)=\widetilde{F}_{n_{0}}(\{(s,1),(s,2)\},z)$, $\alpha_{m_{0_{1}}}=\alpha_{1}$ and similarly for the second state weighting. We present an alternative derivation via a generalised renewal equation and some special cases of Eqs. (29) and (30) in D.
@@
The sustained oscillations, which have only been seen in a small number of systems, including quantum walks [74], biased random walks in hexagonal domains [72] and other non-Markov systems [75] are the natural, higher-dimensional analogue of the bimodality seen in the one-dimensional systems, where again, the peaks correspond to the timescales for which the walker moves away from the target as well as those trajectories that miss the target and must travel round the entire domain. Naturally, as the indirect trajectories take over and more trajectories reach the targets, the oscillations dampen and decay to zero [72].
@@ VI Summary Statistics and the Mean First-Passage Time
The matrices in Eq. (31) are made up of the MFPT between localised sites $\mathdutchcal{F}_{\bm{n}_{0},m_{0}\to\bm{s}_{i},m_{\bm{s}_{i}}}$. More precisely, $\mathbb{T}_{i,i}=0$, $\mathbb{T}_{i,k}=\mathdutchcal{F}_{\bm{s}_{k},m_{\bm{s}_{k}}\to\bm{s}_{i},m_{\bm{s}_{i}}}$ ($i\neq k$), $\mathbb{T}_{{0}_{i,k}}=\mathbb{T}_{i,k}-(2d)^{-1}\sum_{j=1}^{2d}\mathdutchcal{F}_{\bm{n}_{0},m_{0_{j}}\to\bm{s}_{i},m_{\bm{s}_{i}}}$ and $\mathbb{T}_{1_{i,k}}=\mathbb{T}_{i,k}-1$. To find these quantities one evaluates expressions of the form $\mathdutchcal{F}_{\bm{n}_{0},m_{0}\to\bm{s}_{i},m_{\bm{s}_{i}}}=\frac{\partial}{\partial z}\left.\frac{\widetilde{P}_{\bm{n}_{0},m_{0}}(\bm{s},m_{\bm{s}},z)}{\widetilde{P}_{\bm{s},m_{\bm{s}}}(\bm{s},m_{\bm{s}},z)}\right|_{z=1},$ and we give details and explicit quantities in Sec. VIII of ref. [45].
@@
We now make connection with ref. [36], where they consider the global mean first-passage time (GMFPT), $\mathdutchcal{G}_{\{\mathdutchcal{S}\}}$, of a persistent random walk. By performing a summation over $\bm{n}_{0}$ of Eq. (31), and repeatedly using the multi-linearity property of the determinant, we find the GMFPT as
$\mathdutchcal{G}_{\{\mathdutchcal{S}\}}=\frac{\det[\mathbb{G}]}{\det[\mathbb{T}_{1}]-\det[\mathbb{T}]},$ (32)
where $\mathbb{G}_{i,j}=\mathbb{T}_{i,j}-(2dN)^{-1}\sum_{\bm{n}_{0}}\sum_{j=1}^{2d}\mathdutchcal{F}_{\bm{n}_{0},m_{0_{j}}\to\bm{s}_{i},m_{\bm{s}_{i}}}$, while $\mathbb{T}_{1}$ and $\mathbb{T}$ are defined above. In certain situations such as the one target case in periodic domains treated in [36], the GMFPT is independent of the target location, however in general this is not the case. In fact, $\mathdutchcal{G}_{\{\mathdutchcal{S}\}}$ is generally dependent on the location of the target(s), the size of the domain and the movement parameters.
Figure 8: Global mean first-passage time, rescaled by $N$, plotted using Eq. (32). (a) By employing Eq. (S65) in Eq. (32), we show the case of two targets in a one-dimensional reflective domain with one fixed target at $s_{1}=30$ as a function of the placement of the second target $s_{2}$ for varying persistence lengths. (b) Using Eq. (S66) in Eq. (32) we treat the case of an $N\times N$ periodic domain and plot the GMFPT as a function of the rescaled persistence length for different domain sizes. In panel (a) when $s_{1}=s_{2}$ the calculation reduces to a one target case, which explains the slight kink in the plots for this case. In (b) because of the periodicity of the lattice there is no dependence on the target location and we drop the subscript $\{\mathdutchcal{S}\}$ on the $y$-label. To vary $p_{l}$ in this case, we fix $b=0.01$ and vary $f\in[0.02,0.96]$ from low to high values, while simultaneously varying $\ell\in[0.01,0.58]$ from high to low values to ensure normalisation with $c^{(2)}=0.01$. The vertical dashed lines show $p_{l}/N$ values for the diffusive cases.
To analyse this quantity further we consider the quantity of persistence length $p_{l}$, defined in ref. [36] as the mean number of steps in a ballistic excursion. Since our formalisms allow for the inclusion of a sojourn probability, we slightly redefine this quantity as the mean excursion time before a turn takes place. More precisely, this is the mean of a geometric distribution i.e., $p_{l}=\sum_{t=1}^{\infty}t\left[1-\left(f+c^{(\mathfrak{d})}\right)\right]\left(f+c^{(\mathfrak{d})}\right)^{(t-1)}$, which is easily evaluated as $p_{l}=\left[1-\left(f+c^{(\mathfrak{d})}\right)\right]^{-1}$. Using the knowledge that the movement parameters sum up to one (see Sec. IV), one may rewrite this as
$p_{l}=\frac{1}{b+(Z-2)\ell}.$ (33)
In Eq. (33) as $p_{l}$ is defined through the coordination number of the lattice, $Z$, it means that it is valid on all types of lattices. In the diffusive case ($b=\ell=q/Z$) $p_{l}=Z\left[q(Z-1)\right]^{-1}$, where $0<q\leq 1$ is the probability of changing lattice site, with $q=1$ a walker which always moves [6]. In such a case on the square lattice, the movement probabilities are $q/4$ to move in any direction and $1-q$ to remain on the current lattice site, implying $p_{l}=4(3q)^{-1}$.
In Fig. 8, we plot, in panel (a), the GMFPT in a reflecting domain with two targets, with one fixed target at $s_{1}=30$ and varying the position of the second $s_{2}$. In panel (b), we plot a 2D periodic domain with one target and obtain minima as a function of $p_{l}$ as observed also in ref. [36]. To do so, we fix a very low $b=0.01$ and vary $p_{l}$ by changing both $f$ and $\ell$.
In the one-dimensional reflection case, panel (a), an increase in the persistence length lowers the MFPT for all target placement, which is expected due to an increase in persistence always aiding coverage in one-dimension. We also see strong dependence on the placement of the second target with the higher GMFPTs occurring when $s_{1}$ and $s_{2}$ are close together. We note the higher GMFPT when $s_{2}$ is to the right of $s_{1}$. To explain this feature, consider that when $n_{0}<s_{1}<s_{2}$ or $s_{2}<s_{1}<n_{0}$, $s_{2}$ is inaccessible. With the static target at $s_{1}>N/2$, the case where $n_{0}<s_{1}<s_{2}$ leads to, on average, longer MFPTs and also occupies a larger proportion of the contributions to the GMFPT than the $s_{2}<s_{1}<n_{0}$ case. If $s_{1}=N/2$, one would see a symmetric plot around $s_{2}$.